# Actor-Critic Physics-informed Neural Lyapunov Control

### v1

@@ I Introduction
Neural network control policies have shown great promise in model-based control and outperformed existing design methods or even humans, owing to their capability to capture intricate nonlinear policies [1] [2] [3] [4]. For example, neural network control policies trained with reinforcement learning algorithms have been able to outperform human champions in drone racing [4]. Despite the outstanding empirical performance, the application of neural network policies in physical systems is of concern due to a lack of stability and safety guarantees.
Certificate functions such as Lyapunov functions provide a general framework to design nonlinear control policies with provable guarantees. For example, a quadratic Lyapunov function can be found with Linear Quadratic Regulator (LQR) [5] for a linear system, and a polynomial Lyapunov function can be found with Sum-of-Square (SoS) optimization [6] for a polynomial system. However, there is generally no universal way to construct these certificates for general nonlinear systems as it usually requires a predefined function form derived from the structure of the dynamical system.
Along with the development of neural networks as universal function approximators, there have been learning-based methods to construct Lyapunov functions represented by neural networks. The main focus is on either learning a Lyapunov function for an autonomous system [7][8][9] or co-learning a neural network control policy together with a neural Lyapunov function [10] [11] [12]. In these learning-based methods, a typical approach is to minimize the empirical violation of the Lyapunov conditions over the parameters of the controller and/or the certificate. Compared to optimization-based methods such as LQR and SoS, learning-based methods can be applied to more general non-linear systems beyond linear and polynomial systems.
However, in these learning-based methods, the Lyapunov conditions are only guaranteed at the sampled states. Thus there must be a verification procedure after the learning process to certify that the Lyapunov conditions are satisfied over a region. This verification is typically done by using Satisfiability Modulo Theories (SMT) solvers [13]. If the Lyapunov conditions are satisfied over a sub-level set of the neural Lyapunov function, then the learned function is a valid Lyapunov function over the sub-level set and the sub-level set is a domain of attraction (DoA). This brings us to another problem: how to learn a neural network control policy such that the certified region is maximized. Existing learning-based methods usually fail to obtain a large DoA as the loss function is not informed of this objective explicitly, leading to a shape mismatch between the level sets of the learned neural Lyapunov function and that of the true DoA.
@@
In this paper, we propose an algorithmic framework for learning a neural network control policy and a neural network Lyapunov function for actuation-constrained nonlinear systems. Our starting point is to use a physics-informed loss function based on Zubov’s Partial Differential Equation (PDE), the solution of which (the Zubov function) characterizes the true DoA. The framework follows an actor-critic pattern where we switch between learning a Zubov function and improving the control policy by minimizing its time derivative. By minimizing the PDE residual, the Our numerical experiments on several stabilization tasks show that our approach can enlarge the DoA significantly compared to the state-of-the-art approaches. The source code is available at https://github.com/bstars/ZubovControl
@@ I-A Related Work
There is a large body of work on learning Lyapunov functions for autonomous systems [7] [8] [9] [14]. In the work of Abate et al [7], they proposed a counter-example guided framework to learn a neural Lyapunov function. The algorithm switches between minimizing the violation of Lyapunov conditions at some sampled points and using an SMT solver to either verify the Lyapunov conditions or falsify them by counter-examples that are added to the training set. A common feature of these methods is that the level set of the learned Lyapunov function does not match the true DoA, leading to a conservative estimation of the DoA despite using neural networks. To capture the maximum region of attraction, Liu et al [15] mapped an easily constructed Lyapunov function, the time integration of the norm of the state along a trajectory, which is a true Lyapunov function in nature, to a Zubov function and learned a neural Zubov function with a physics-informed neural network. The result showed that this method approximately captures the maximum region of attraction. Our method also utilizes Zubov’s equation to co-learn the control policy and the stability certificate.
@@
This work is also built upon existing works on co-learning Lyapunov functions and control policies. Chang et al [10] proposed a counter-example guided method similar to [7] to co-learn a Lyapunov function and a control policy that jointly satisfy the Lyapunov conditions for nonlinear systems. In their work, they add a regularizer to the loss function to enlarge the DoA. The method in [10] has also been extended to safety certificates [11], unknown nonlinear systems [12] and discrete-time systems [16]. In our method, a regularizer is unnecessary as the Zubov equation itself captures the true DoA.
@@
Actor-critic methods can be seen as an approximation of policy iteration [18]. There have also been works that combine neural Lyapunov functions with policy iteration to learn formally verifiable control policies. In the work of [19], they use policy iteration to learn a control policy for a control-affine system. The policy evaluation part is done by representing the value function by a physics-informed neural network and minimizing the violation of the Generalized Hamilton-Jacobi-Bellman equation [20], and the policy improvement is done by minimizing the value function. Our method also follows a policy-evaluation-policy-improvement paradigm as in actor-critic method and policy iteration.
[…]
@@ II-A Lyapunov Certificates
where $x\in\mathcal{D}\subseteq\mathbb{R}^{n}$ is the state and $f\colon\mathcal{D}\to\mathbb{R}^{n}$ is a Lipschitz continuous function. For an initial condition $x_{0}\in\mathcal{D}$, we denote the unique solution by $x(t;x_{0})$, which we assume exists for all $t\geq 0$. Without loss of generality, we assume that the origin is an equilibrium of the system, $f(0)=0$.
@@
The zero solution $x(t)=0$ to (1) is (locally) asymptotically stable if there exists a $\delta>0$ such that if $\|x_{0}\|<\delta$, then $\lim_{t\to\infty}x(t;x_{0})=0$
@@
If the solution $x(t)=0$ to (1) is asymptotically stable, then the domain of attraction (DoA) is given by
$\mathcal{A}:=\big{\{}x_{0}\in\mathcal{D}:\lim_{t\to\infty}\|x(t;x_{0})\|=0\big{\}}$
@@
Theorem 1 (Lyapunov Stability)
Consider the nonlinear dynamical system in (1). If there exists a continuously differentiable function $V\colon\mathcal{D}\rightarrow\mathbb{R}$ satisfying the following conditions
@@
then $x(t)=0$ is asymptotically stable. Furthermore, any sub-level set $\mathcal{D}_{\beta}=\{x:V(x)\leq\beta\}\subseteq\mathcal{D}$ containing the origin is a subset of the domain of attraction [21].
To find the maximum estimate of the domain of attraction, one must find the maximum $\beta$ such that $\mathcal{D}_{\beta}\subseteq\mathcal{D}$.
@@ II-B Problem Statement
where $x\in\mathcal{D}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}$ is the control input, $\mathcal{U}\subseteq\mathbb{R}^{m}$ is the actuation constraint set, and $f\colon\mathcal{D}\times\mathcal{U}\to\mathbb{R}^{n}$ is a Lipschitz continuous function. In this paper, we assume that $\mathcal{U}$ is a convex polyhedron.
Given a locally-Lipschitz control policy $\pi(\cdot)$ and an initial condition $x_{0}\in\mathbb{R}^{n}$, we denote the solution of the closed-loop system $\dot{x}=f(x,\pi(x))=f_{cl}(x)$ at time $t\geq 0$ by $x(t,x_{0};\pi)$.
In this paper, we are interested in learning a nonlinear control policy $\pi_{\gamma}(\cdot)$, parameterized by a neural network with trainable parameters $\gamma$, with the following desiderata:
@@
    The domain of attraction of the equilibrium is maximized.
@@ II-C Neural Lyapunov Control
A typical learning-enabled approach to designing stabilizing controllers is to parameterize the Lyapunov function candidate $V_{\theta}$ and the controller $\pi_{\gamma}$ with neural networks and aim to minimize the expected violation of Lyapunov conditions (1), i.e.,
$\min_{\theta,\gamma}\!E_{x\sim\rho}\!\big{[}V_{\theta}(0)^{2}\!+\!(-V_{\theta}(x))_{+}\!+\!\big{(}\nabla_{x}V_{\theta}(x)^{\top}f(x,\pi_{\gamma}(x))\big{)}_{+}\big{]}$ (4)
where $\rho$ is a predefined sampling distribution with support on $\mathcal{D}$ [10].
There are potentially infinitely many Lyapunov function candidates that minimize the expected loss, including the shortcut solution $V_{\theta}(x)=0~{}\forall x$. To avoid this shortcut and promote large domains of attractions, we must use regularization, e.g., $(\|x\|_{2}-\alpha V_{\theta}(x))^{2}$ [10]. However, poorly chosen regularizers can lead to a mismatch between the shape of level sets of the learned Lyapunov function and the true DoA. Motivated by this drawback, we will design a physics-informed loss function to directly capture the true DoA.
@@ III-A Maximal Lyapunov Functions and Zubov’s Method
To characterize the domain of attraction exactly, we can use a variation of the Lyapunov function, called the maximal Lyapunov function [22].
Theorem 2 (Maximal Lyapunov Function)
@@
Then $\mathcal{A}$ is the domain of attraction.
@@
Due to the presence of infinity, the maximal Lyapunov function is hard to represent by function approximators such as neural networks. Another way to build a Lyapunov function and characterize the domain of attraction is using Zubov’s Theorem [21] [23].
@@
Consider the nonlinear dynamical system in (1) with $f(0)=0$. Let $\mathcal{A}\subseteq\mathcal{D}$ and assume there exists a continuously differentiable function $W\colon\mathcal{A}\rightarrow\mathbb{R}$ and a positive definite function Ψ:nnormal-:normal-$\Psi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfying the following conditions,
@@
$W(x)\to 1\text{ as }x\to\partial\mathcal{A}$ (7c)
@@
Then the zero solution $x(t)=0$ is asymptotically stable with domain of attraction $\mathcal{A}$.
@@
We note that the $V(\cdot)$ function in Theorem 2 and the $W(\cdot)$ function in Theorem 3 can be related by [24]
@@ III-B Learning the Zubov Function (Critic)
Suppose we have a fixed policy $\pi_{\gamma}(\cdot)$ and the corresponding closed-loop system $\dot{x}=f(x,\pi_{\gamma}(x))$. Let $V(\cdot)$ be chosen as in (6). The relationship (8a), combined with (7d), enables us to learn the neural Zubov function $W_{\theta}(x)\approx\mathrm{tanh}(\alpha V(x))$, which can be interpreted as an evaluation of the policy $\pi_{\gamma}$. To this end, we use the following loss function
@@
$L_{p}(\theta)$ $=\frac{1}{K}\sum_{i=1}^{K}\bigg{[}\nabla_{x}W_{\theta}(x_{i})^{\top}f(x_{i},\pi(x_{i}))+$ (10c)
$\qquad\qquad\alpha(1-W_{\theta}(x_{i}))(1+W_{\theta}(x_{i}))\Phi(x)\bigg{]}^{2}$
Here $W_{\theta}(\cdot)$ is a neural network with $\mathrm{tanh}$ activation functions at the output layer [25] such that $W_{\theta}(x)\in(0,1)\forall x\in\mathcal{D}$. The first term $L_{z}(\theta)$ penalizes the violation of the condition (7a). The second term $L_{r}(\theta)$ penalizes the violation of the relationship (8a). To evaluate $V(\cdot)$ that appears in this loss, we must simulate the closed-loop system from the initial conditions $\{x_{i}\}_{i=1}^{M}$. Finally, the third term $L_{p}(\theta)$ is the physic-informed part to minimize the residual of the Zubov PDE in (7d).
@@ III-C Policy Improvement (Actor)
Given a policy $\pi_{\gamma}(\cdot)$, the corresponding Zubov function is $W(x)=\tanh(\alpha V^{\pi_{\gamma}}(x))$ where $V^{\pi_{\gamma}}(x)$ is defined as in (6). We can improve the policy to drive the state to $x=0$ by minimizing the time derivative $\dot{W}(x)$. However, since we cannot differentiate $W(x)$ with respect to $x$, we use $W_{\theta}(\cdot)$ as a proxy of the true $W$, which results in the following loss function
@@ III-D Actor-Critic Learning
[…]
@@
This loss function stabilizes the training at the beginning.
@@
where $M$ is the number of trajectories sampled at each iteration, $K$ is the batch size. Note that rather than alternating between updating $W_{\theta}(\cdot)$ and $\pi_{\gamma}(\cdot)$ with the other one fixed, are use the $\mathbf{sg(\cdot)}$ operator to enable a simultaneous update of $W_{\theta}(\cdot)$ and $\pi_{\gamma}(\cdot)$. We outline the overall method in 1.

### v2

@@ I Introduction
Neural network control policies have shown great promise in model-based control and outperformed existing design methods, owing to their capability to capture intricate nonlinear policies [1] [2] [3] [4]. For example, neural network control policies trained with reinforcement learning algorithms have been able to outperform human champions in drone racing [4]. Despite the outstanding empirical performance, the application of neural network policies in physical systems is of concern due to a lack of stability and safety guarantees.
Certificate functions such as Lyapunov functions provide a general framework for designing nonlinear control policies with provable guarantees. For instance, a quadratic Lyapunov function can be derived using a Linear Quadratic Regulator (LQR) [5] for linear systems, while a polynomial Lyapunov function can be obtained through Sum-of-Squares (SoS) optimization [6] for polynomial systems. However, as systems become more complex and nonlinear, the automated construction of Lyapunov functions becomes an increasingly crucial research focus within control theory [7, 8, 9, 10].
Learning-based methods combined with neural networks have been a promising alternative to optimization-based methods in constructing Lyapunov certificates for autonomous systems [8][11][12] or co-learning a neural network control policy together with a neural Lyapunov function [13, 14, 15]. In these learning-based methods, a typical approach is to minimize the violation of the Lyapunov conditions at a finite number of sampled states, and then invoke a verifier such as SMT solvers [16] to extend the validity of the learned certificate beyond the sampled states. Although applicable to more general nonlinear systems, the choice of the training loss function is critical to the learner’s success. For example, for the case of stability analysis, we are interested in certifying the largest region of attraction, which is missing in a training objective that merely penalizes the Lyapunov condition violations.
@@
In this paper, we propose an algorithmic framework for co-learning a neural network control policy and a neural network Lyapunov function for actuation-constrained nonlinear systems. Our starting point is to use a physics-informed loss function based on Zubov’s Partial Differential Equation (PDE), the solution of which characterizes the ground domain of attraction (DoA) for a given control policy. Our framework then follows an actor-critic pattern where we alternate between learning a Zubov function, by minimizing the PDE residual, and improving the control policy, by minimizing its Lie derivative akin to Sontag’s formula [17]. To improve the efficiency of the method, we provide an implementation that essentially runs the actor and critic updates in parallel. Finally, we propose a novel method based on the softmax function to enforce polytopic actuation constraints without any projection layer. Our numerical experiments on several stabilization tasks show that our approach can enlarge the DoA significantly compared to state-of-the-art techniques. The source code is available at https://github.com/bstars/ZubovControl.
@@ I-A Related Work
There is a large body of work on learning Lyapunov functions for autonomous systems [8] [11] [12] [18]. In the work of Abate et al [8], they proposed a counter-example guided framework to learn a neural Lyapunov function. A common feature of these methods is that the level set of the learned Lyapunov function does not match the true DoA, leading to a conservative estimation of the DoA despite using neural networks. Liu et al [19] used Zubov’s PDE as a training loss to capture the maximum region of attraction for autonomous systems. Our method also utilizes Zubov’s equation to co-learn the control policy and the stability certificate together.
@@
Chang et al. [13] proposed a counter-example guided method similar to [8] to co-learn a Lyapunov function and a control policy that jointly satisfy the Lyapunov conditions for nonlinear systems. In their work, they add a regularizer to the loss function to enlarge the DoA. The method in [13] has also been extended to safety certificates [14], unknown nonlinear systems [15] and discrete-time systems [20]. In our method, regularization is unnecessary as the Zubov equation captures the true DoA.
@@
Actor-critic methods can be seen as an approximation of policy iteration [22]. There have also been works that combine neural Lyapunov functions with policy iteration to learn formally verifiable control policies. In the work of [23], they use policy iteration to learn control policies for control-affine systems. The policy evaluation part is done by representing the value function by a physics-informed neural network and minimizing the violation of the Generalized Hamilton-Jacobi-Bellman equation [24], and the policy improvement is done by minimizing the value function. Our method also follows a policy-evaluation-policy-improvement paradigm as in the actor-critic method and policy iteration.
The rest of the paper is organized as follows. In§$\S$§II, we provide background and problem statement. In §$\S$§III, we describe how to co-learn a control policy and a Lyapunov function with a physics-informed neural network and the verification procedure after training. We provide numerical experiments in §$\S$§IV.
@@ II-A Lyapunov Certificates
where $x\in\mathcal{D}\subseteq\mathbb{R}^{n}$ is the state and $f\colon\mathcal{D}\to\mathbb{R}^{n}$ is a Lipschitz continuous function. For an initial condition $x_{0}\in\mathcal{D}$, we denote the unique solution by $x(t;x_{0})$, which we assume exists for all $t\geq 0$. We assume that the origin is an equilibrium of the system, $f(0)=0$.
Definition 1 (Stability[25])
The zero solution $x(t)=0$ to (1) is stable if for any $\epsilon>0$, there exists $\delta>0$ such that if $\|x_{0}\|<\delta$, then $\|x(t;x_{0})\|<\epsilon$ for all $t>0$.
@@
The zero solution $x(t)=0$ to (1) is locally asymptotically stable if it is stable and there exists a $\delta>0$ such that if $\|x_{0}\|<\delta$, then $\lim_{t\to\infty}\|x(t;x_{0})\|=0$.
@@
If the zero solution $x(t)=0$ to (1) is asymptotically stable, then the domain of attraction (DoA) is given by $\mathcal{A}:=\big{\{}x_{0}\in\mathcal{D}:\lim_{t\to\infty}\|x(t;x_{0})\|=0\big{\}}$.
@@
Theorem 1 (Lyapunov Stability[25])
Consider the dynamical system in (1). If there exists a continuously differentiable function $V\colon\mathcal{D}\rightarrow\mathbb{R}$ satisfying the following conditions
@@
then the zero solution $x(t)=0$ is asymptotically stable.
Any sub-level set $\mathcal{D}_{\beta}=\{x:V(x)\leq\beta\}\subseteq\mathcal{D}$ containing the origin is a subset of DoA [25], and the largest $\mathcal{D}_{\beta}$ can be calculated by maximizing $\beta$ subject to $\mathcal{D}_{\beta}\subseteq\mathcal{D}$.
@@ II-B Problem Statement
where $x\in\mathcal{D}\subseteq\mathbb{R}^{n}$ is the state, $u\in\mathcal{U}$ is the control input, $\mathcal{U}\subseteq\mathbb{R}^{m}$ is the actuation constraint set, and $f\colon\mathcal{D}\times\mathcal{U}\to\mathbb{R}^{n}$ is a Lipschitz continuous function. This paper assumes that $\mathcal{U}$ is a convex polyhedron.
Given a locally-Lipschitz control policy $\pi(\cdot)$ and an initial condition $x_{0}\in\mathbb{R}^{n}$, we denote the solution of the closed-loop system $\dot{x}=f(x,\pi(x))=f_{cl}(x)$ at time $t\geq 0$ by $x(t,x_{0};\pi)$. In this paper, we are interested in learning a nonlinear control policy $\pi_{\gamma}(\cdot)$, parameterized by a neural network with trainable parameters $\gamma$, with the following desiderata:
@@
    The DoA of the zero solution is maximized.
@@ II-C Neural Lyapunov Control
A typical learning-enabled approach to designing stabilizing controllers is to parameterize the Lyapunov function candidate $V_{\theta}$ and the controller $\pi_{\gamma}$ with neural networks and aim to minimize the expected violation of Lyapunov conditions (2)
$\min_{\theta,\gamma}\!E_{x\sim\rho}\!\big{[}V_{\theta}(0)^{2}\!+\!(-V_{\theta}(x))_{+}\!+\!\big{(}\nabla_{x}V_{\theta}(x)^{\top}f(x,\pi_{\gamma}(x))\big{)}_{+},\big{]}$ (4)
where $\rho$ is a predefined sampling distribution with support on $\mathcal{D}$ [13]. There are potentially infinitely many Lyapunov function candidates that minimize the expected loss, including the shortcut solution $V_{\theta}(x)=0~{}\forall x\in\mathcal{D}$. To avoid this shortcut and promote large DoA, we must use regularization, e.g., $(\|x\|_{2}-\alpha V_{\theta}(x))^{2}$ [13]. However, poorly chosen regularizers can lead to a mismatch between the shape of level sets of the learned Lyapunov function and the true DoA. Motivated by this drawback, we will incorporate a physics-informed loss function to directly capture the true DoA.
@@ III-A Maximal Lyapunov Functions and Zubov’s Method
To characterize the DoA exactly, we can use a variation of the Lyapunov function, the maximal Lyapunov function [26].
Theorem 2 (Maximal Lyapunov Function[26])
@@
Then $\mathcal{A}$ is the DoA for the system in (1).
@@
Due to the presence of infinity in the integral, the above maximal Lyapunov function is hard to represent by function approximators such as neural networks. Another way to build a Lyapunov function and characterize the DoA is using Zubov’s Theorem [25] [27].
@@
Consider the nonlinear dynamical system in (1). Let $\mathcal{A}\subseteq\mathcal{D}$ and assume there exists a continuously differentiable function $W\colon\mathcal{A}\rightarrow\mathbb{R}$ and a positive definite function $\Psi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfying the following conditions,
@@
$W(x)\to 1\text{ as }x\to\partial\mathcal{A}\text{ or }\|x\|\to\infty$ (7c)
@@
Then $x(t)=0$ is asymptotically stable with DoA $\mathcal{A}$.
@@
The maximal Lyapunov function $V(\cdot)$ of Theorem 2 and the Zubov function $W(\cdot)$ of Theorem 3 can be related by [28]
@@ III-B Learning the Zubov Function (Critic)
Suppose we have a fixed policy $\pi_{\gamma}(\cdot)$ and the corresponding closed-loop system $\dot{x}=f(x,\pi_{\gamma}(x))$. Let $V(\cdot)$ be chosen as in (6). The relationship (8a), combined with (7d), enables us to learn the Zubov function $W_{\theta}(x)\approx\mathrm{tanh}(\alpha V(x))$, which can be interpreted as an evaluation of the policy $\pi_{\gamma}$. Here $W_{\theta}(\cdot)$ is a neural network with $\mathrm{sigmoid}$ activation functions at the output layer [29] such that $W_{\theta}(x)\in(0,1)\ \forall x\in\mathcal{D}$. To train $W_{\theta}(\cdot)$, we use the following loss function,
@@
$L_{p}(\theta)$ $=\frac{1}{K}\sum_{i=1}^{K}\bigg{[}\nabla_{x}W_{\theta}(x_{i})^{\top}f(x_{i},\pi_{\gamma}(x_{i}))+$ (10c)
$\qquad\qquad\alpha(1-W_{\theta}(x_{i}))(1+W_{\theta}(x_{i}))\Phi(x_{i})\bigg{]}^{2}$
The first and second terms, $L_{z}(\theta)$ and $L_{r}(\theta)$, penalize the violation of (7a) and (8a), respectively. We note that to evaluate $V(\cdot)$ that appears in this loss, we simulate the closed-loop system from $M$ initial conditions $\{x_{i}\}_{i=1}^{M}$. Finally, the third term $L_{p}(\theta)$ is the physic-informed part to minimize the residual of the Zubov PDE in (7d) at $K$ sampled states.
@@ III-C Policy Improvement (Actor)
Given a policy $\pi_{\gamma}(\cdot)$, the corresponding gound-truth Zubov function is $W(x)=\tanh(\alpha V^{\pi_{\gamma}}(x))$ where $V^{\pi_{\gamma}}(x)$ is defined as in (6) for the autonomous system $\dot{x}=f(x,\pi_{\gamma}(x))$. We can improve the policy to drive the state to $x=0$ by minimizing the Lie derivative of $W$ along $f$ akin to Sontag’s formula [17]. However, since we cannot differentiate $W(x)$ with respect to $x$, we use $W_{\theta}(\cdot)$ as a proxy of the true $W$, which results in the following loss function
@@ III-D Actor-Critic Learning
We now combine the critic of §$\S$§III-B and the actor of §$\S$§III-C to co-learn the controller and the Zubov function. We first define $R_{1}\subseteq\mathcal{D}$ to be the region from which we sample $\{x_{i}\}$ and define the region $R_{2}=\{ax:x\in R_{1}\}$ ($a>1$) as the region of interest. In this paper we choose $a=2$. Since the controller is randomly initialized, the sampled trajectories might diverge. To prevent this, we add a loss function as follows,
@@
Since $W(\cdot)$ is constrained to be in $[0,1]$ and the actor is trained to minimize the Lie derivative of ${W}$, this loss prevents the states from going to $\partial R_{2}$, where $W(\cdot)$ takes the maximum value $1$. This loss function stabilizes the training at the beginning.
@@
Note that rather than alternating between updating $W_{\theta}(\cdot)$ and $\pi_{\gamma}(\cdot)$ with the other one fixed, we use the $\mathbf{sg(\cdot)}$ operator to enable a simultaneous update of $W_{\theta}(\cdot)$ and $\pi_{\gamma}(\cdot)$. We outline the overall method in Algorithm 1.